Bon has 18 coins. If he has $3.45 and only quarters and dimes, how many of each does he have?
Answer by checkley77(12844)

Q+D=18
Q=18-D
.25Q+.10D=3.45
.25(18-D)+.10D=3.45
4.50-.25D+.10D=3.45
-.15D=3.45-4.50
-.15D=-1.05
D=-1.05/-.15
D=7 DIMES.
Q+7=18
Q=18-7
Q=11 QUARTERS.
PROOF:
.25*11+.10*7=3.45
2.75+.70=3.45
3.45=3.45